# Concept Formation and Alignment in Language Models: Bridging Statistical Patterns in Latent Space to Concept Taxonomy

### v1

@@ Abstract
This paper explores the concept formation and alignment within the realm of language models (LMs). We propose a mechanism for identifying concepts and their hierarchical organization within the semantic representations learned by various LMs, encompassing a spectrum from early models like Glove to the transformer-based language models like ALBERT and T5. Our approach leverages the inherent structure present in the semantic embeddings generated by these models to extract a taxonomy of concepts and their hierarchical relationships. This investigation sheds light on how LMs develop conceptual understanding and opens doors to further research to improve their ability to reason and leverage real-world knowledge. We further conducted experiments and observed the possibility of isolating these extracted conceptual representations from the reasoning modules of the transformer-based LMs. The observed concept formation along with the isolation of conceptual representations from the reasoning modules can enable targeted token engineering to open the door for potential applications in knowledge transfer, explainable AI, and the development of more modular and conceptually grounded language models.
@@ 1 Introduction
Language models are rapidly approaching human-level language abilities. However, these advancements raise important concerns - on the one hand, despite their immense potential, language models are susceptible to unexpected and potentially harmful behaviors such as hallucination [21], stereotyping [25], misinformation [37], and leakage of sensitive training data [29]. On the other hand, the extent of their capabilities and the nature of their understanding is not yet known, thus capturing the attention of diverse groups, from policymakers to academics [16, 39, 3, 18]. While these issues may seem different, they both stem from a common concern - the "black box" nature. The limited ability to understand how they reach their outputs raises fears about their true capabilities, potential biases, and unintended consequences.
Researchers are actively studying the semantic representations learned by large language models (LLMs), aiming to understand the inner workings of language models. While initially used to study the learning dynamics of linguistic features and abstractions within LMs [50, 26], this research has expanded to explore how these models acquire and represent knowledge about the world. Some studies [38, 18, 1] have explored the ability of transformer-based models to learn representations of color, spatial, and temporal information. In contrast, other studies have raised concerns about the coherence and completeness of this knowledge [16]. The work of [39] even suggests that language models may lack a true conceptual understanding of the real world, echoing the "stochastic parrot" critique [3]. This debate stems from the fundamental question of whether language models merely capture statistical patterns in language or truly grasp the underlying concepts they represent. There are so many works on the semantic relations of the tokens but limited research directly addresses the question of whether these relations are grounded in a conceptual understanding of the real world, or merely reflect co-occurrence patterns in language data [2, 28].
Objectives. To address the aforementioned gap, this paper investigates the process of concept formation in language models. We posit our work based on the argument that ‘Concepts’ are mental representations of categories, which group objects, events, or ideas based on shared features [45]. More generic concepts that correspond to broader categories and share fewer defining features. This is how the hierarchical categories/concepts are formed. In an unsupervised world, when there is no label, categories translate into clusters/communities. Then, similar to [28], we argue that if neural models of language are to have any knowledge about concepts, it ought to be reflected in their token/word embeddings. Thus, communities/categories need to form in the semantic space in which their most relevant category names are often closely related to concepts that describe that category. Note that, while we often form concepts based on categories, not every category will necessarily lead to a clear and distinct concept in the real world. This phenomenon, known as concept alignment [43] will be the focus of our quantitative analysis.
Methodology. Interpretability research traditionally use probing [36, 51]. Probing often involves evaluating models on specific, often artificial tasks designed to isolate particular linguistic phenomena. Probing results might be sensitive to the specific design choices of the probe tasks, such as the choice of dataset, prompts, or metrics, making it difficult to extrapolate conclusions to the model’s overall capabilities in real-world scenarios. Thus, while effective for specific domains, they have inherently limited ability to provide broader insights into the extent of models’ underlying representations and reasoning capabilities beyond specific tasks they are tuned for [48] Therefore, we avoid using conventional probing, instead, we propose a concept extraction methodology based on graph creation (K-nearest neighbors), fuzzy weighting mechanism (borrowed from UMAP category theory [32]) and community/cluster detection algorithms to identify groups of highly interconnected data points, representing potential concepts. The performance of this approach hinges upon the existence of a meaningful, hierarchical structure within the input embedding layer. This method can potentially offer a deeper and more holistic understanding of the underlying knowledge encoded in these representations (as compared to probing).
Models Under Investigation. Using our proposed method, we investigate the embedding spaces of GloVe, a traditional statistical language model [40]; ALBERT, a transformer-based encoder language model [23]; and T5, a transformer-based autoregressive language model [42] to comprehend if concept formation occurs in any of these models.
Evaluation. In order to answer the question of "if concept formation happens in the semantic space," we evaluated the formed clusters against external references such as WordNet [34], Name Dataset [44], and country-state-city databases [11] for most of our evaluations. For clusters where open-source datasets were not available, we employed human annotators as an external reference. We observed that the process of concept formation can occur as early as the input embedding layer in transformer-based models. Considering that this observation hints at the existence of semantic memory in transformer-based models, we further investigate in Section 4.3 whether there is a separation between the identified semantic memory and the reasoning mechanisms. Specifically, we ask whether the model can still reason effectively with a modified semantic memory. This line of inquiry aims to deepen our understanding of the relationship between knowledge representation and reasoning capabilities within transformer-based models.
Figure 1: Simplified steps on how external information is understood and retained. Upon understanding a newly encountered word/entity, it is typically stored in the semantic memory. The existence of semantic memory (on the left) allows the previously encountered words/entities to have a form of meaning even without requiring an external context. The scatter box on the right is the community (primarily associated with moving creatures) we extracted from the ALBERT model [23].
2 Semantic Representation Space
2.1 Concept formation, Representation Learning and Semantic Memory
A simplified model of human understanding can be described as a multi-step process in which the incoming sensory information is integrated and turned into a brain-constructed interpretation of external objects/stimuli. This results in a mental construct (or percept) [15]. Then, the newly formed mental construct gets compared, integrated, and associated with the existing knowledge already retained in the semantic memory. These connections help us to conceptualize and understand the new information based on what we already know (as shown in Figure 1). While diverse reasoning mechanisms exist, a common thread among them seems to be their reliance on and interaction with the semantic memory. This interaction likely leaves enduring traces within such a memory [5]. The integration of newly encountered mental constructs within an existing cognitive framework is often guided by their relations to established internal constructs [13]. This process facilitates the creation, refinement, or expansion of semantically similar clusters [27], implying a degree of inherent categorization [^1] This notion implies an interrelation between the recognition and categorization..
Representation Learning and Concept Formation Representation learning is a fundamental aspect of language models, where the goal is to learn distributed representations (embeddings) for words or subword units that capture their semantic and syntactic relationships [4]. Early models like word2vec [33] utilized shallow neural networks to learn these embeddings from large text corpora. More recent models, such as Large Language Models (LLMs) like GPT-3 [30], employ transformer architectures with self-attention mechanisms [53], enabling more accurate and dynamic representation learning. From the distributional hypothesis, models can form "concepts" by identifying patterns and relationships within data, particularly through recognizing the approximate invariance of shared features across different data instances [13]. It should be noted that "concept formation" is a byproduct of the learning process, not a guaranteed outcome. Furthermore, the formed "concepts" might not always align with human-defined concepts. Thus, it is crucial to investigate and evaluate the nature of concept formation within models and their alignment with human understanding.
2.2 Evaluation Methods
Intrinsic evaluation methods, assess semantic representations by measuring word embedding similarity, comparing them to human judgments of relatedness, (e.g. [33, 17, 52, 36, 51]). While these methods provide valuable insights, they often suffer from the limitations that probing techniques impose, and may not fully capture how well a model forms conceptual hierarchies or grasps the full spectrum of semantic relationships, as word embeddings typically represent individual words rather than higher-level categories [9]. This limitation can hinder the evaluation of a model’s ability to understand broader semantic relationships and its capacity for abstract reasoning.
Another approach is to use extrinsic evaluation methods, which measure the performance of language models on downstream tasks that rely on semantic understanding, such as sentiment analysis, question answering, and machine translation [54, 55]. The performance on these tasks can indirectly indicate the quality of the learned semantic representations. The main challenge with using such extrinsic evaluation metrics to assess concept formation in the embedding space is that they do not directly measure the quality of the concepts themselves. Instead, they measure how well the model performs on specific tasks that rely on those concepts, which does not necessarily prove that the model has formed robust, human-like concepts. For instance, a model might accurately classify sentiment without truly grasping the nuances of emotions like sarcasm or irony [31].
In addition to evaluation metrics, researchers have also employed visualization methods to interpret learned representations. These methods often involve projecting high-dimensional embeddings into lower-dimensional spaces for visualization [48, 46]. However, several challenges arise with these approaches. Firstly, they offer indirect and subjective assessments of model understanding, lacking a quantitative basis for evaluation. Secondly, the dimensionality reduction process can lead to information loss [14, 56], potentially obscuring important nuances in the representations. Finally, different visualization techniques can produce conflicting results, making it difficult to reach definitive conclusions about the model’s true comprehension. To mitigate these shortcomings, we propose a concept extraction mechanism that identifies communities in higher-dimensional space. This approach enables both quantitative evaluation and mitigation of potential information loss incurred during dimensionality reduction.
@@ 3 Concept Extraction
As discussed in Section 2.1, identifying clusters or communities within the semantic representation space is a crucial initial step in evaluating concept formation. We prioritize community detection among clustering methods for the following reasons:
  • Alignment with Semantic Network Structure: The underlying semantic representation, as detailed in Section 2.1, exhibits a network-like structure. Community detection algorithms are specifically designed to uncover such inherent community structures within networks, making them better suited to our task.
  • Unknown Number of Clusters: Traditional clustering algorithms are often sensitive to the choice of the number of clusters or other clustering parameters [57]. In our case, we lack this prior knowledge and only assume that semantically similar words are closer in the representation space. Community detection methods do not necessitate a predetermined number of clusters.
  • High Dimensionality of Data: The high dimensionality of the semantic representation space poses challenges for many clustering algorithms [57]. Community detection methods on the other hand, naturally circumvent this issue by transforming the high-dimensional space into a graph , where relationships between data points become more readily apparent.
This approach allows for a more effective and accurate identification of semantic clusters, reflecting the natural structure of the semantic representation space and facilitating a deeper understanding of concept formation.
@@ 3.1 Graph Construction
While categories themselves are not strict clusters, they may exhibit internal structure or fuzzy partitions based on shared features or relationships. Considering our discussion in the previous section, we seek a representation space where the structure of these categories is explicitly reflected in the geometry of the manifold. This requirement arises from the desire to directly manipulate and modify these clusters through geometric operations on the representation, opening up possibilities for Token Engineering within the language models.
Inspired by the effectiveness of the UMAP dimensionality reduction [32], we use the Nearest neighbor descent to find the $K$ nearest neighbors for every token embedding in the embedding space and then use the following equations as the weight function of the edge between $x_{i}$ and $x_{j}$ nodes in the $K$-NN graph [32]:
@@
Building upon our theoretical arguments, the invariant representations (if they exist), should form fuzzy partitions that can be detected by the community detection algorithms. Thus, for the third step, we perform community detection.
@@ 3.2 Louvain Community Detection
The Louvain method is a widely used algorithm for community detection in large networks. It finds the communities by optimizing a metric called modularity. The modularity of a partition is a scalar value between -1 and 1 that measures how much more densely connected the nodes within a community are compared to how connected they would be in a random network. [6]. For a weighted graph, modularity is defined as:

### v3

# Aligned at the Start: Conceptual Groupings in LLM Embeddings

@@ Abstract
This paper shifts focus to the often-overlooked input embeddings – the initial representations fed into transformer blocks. Using fuzzy graph, k-nearest neighbor (k-NN), and community detection, we analyze embeddings from diverse LLMs, finding significant categorical community structure aligned with predefined concepts and categories aligned with humans. We observe these groupings exhibit within-cluster organization (such as hierarchies, topological ordering, etc.), hypothesizing a fundamental structure that precedes contextual processing. To further investigate the conceptual nature of these groupings, we explore cross-model alignments across different LLM categories within their input embeddings, observing a medium to high degree of alignment. Furthermore, provide evidence that manipulating these groupings can play a functional role in mitigating ethnicity bias in LLM tasks.
Aligned at the Start: Conceptual Groupings in LLM Embeddings
Mehrdad Khatir Virginia Tech khatir@vt.edu                        Sanchit Kabra Virginia Tech sanchit23@vt.edu                        Chandan K. Reddy Virginia Tech reddy@cs.vt.edu
@@ 1 Introduction
Large Language models (LLMs) are rapidly approaching human-level language abilities. However, these advancements raise important concerns - despite their immense potential, language models are susceptible to unexpected and potentially harmful behaviors such as hallucination 21, stereotyping 26, misinformation 36, and leakage of sensitive training data 29. Moreover, the extent of their capabilities and the nature of their understanding is not yet known, thus capturing the attention of diverse groups, from policymakers to academics 17, 39, 3, 19. While these issues may seem orthogonal, they stem from a common concern - the "black box" nature of large language models. The limited ability to understand how they reach their outputs raises doubts about their true capabilities, potential biases, and unintended consequences.
A crucial step towards addressing these challenges lies in understanding how LLMs internally represent and manipulate semantic concepts, the building blocks of human language and thought. Unraveling these internal representations is essential for not only improving LLM performance and robustness, but also for building trustworthy machine learning systems.
In the realm of contextual representations, some studies 38, 19, 2 have investigated the ability of transformer-based models to learn representations of color, spatial, and temporal information. These studies often rely on analyzing contextual embeddings, which are intermediate outputs of LLMs. Moreover, contextual embeddings are inherently tied to the specific input context, making them highly variable, less generalizable, and less controllable/manipulatable.
In our work, we explore this critical area by investigating the emergence of conceptual structures within the input embedding space of LLMs. We aim to explore if these conceptual formations form independent of the context, and if such groupings are in alignment across LLMs, if these clusters exhibit internal organization, and if these patterns have functional implications for model behavior. Specifically, we examine:
  • RQ1: Whether semantically related words and phrases are grouped together, forming identifiable conceptual clusters that are aligned with external world concepts and categories.
  • RQ2: Do these groupings exhibit intra-cluster organization such as hierarchies, topological ordering, etc., thereby suggesting the formation of structured concepts.
  • RQ3: Is there an inter-model alignment in semantic organization, across diverse transformer-based LLMs (Albert 23, T5 42, Llama3 1) irrespective of architecture, size, or pretraining data.
  • RQ4: Whether these groupings play a functional role in LLMs. We test this by a case study that tries to mitigate ethnicity bias through cluster modification.
To uncover conceptual structures within the input embedding space of LLMs, we employ fuzzy graph construction 32. Then, the fuzzy graph is analyzed using a community detection algorithm 6 to reveal conceptual groupings and their categorical organization. We use Louvain community detection, which is effective at revealing hierarchical community structures 6 in conjunction with multiple k choices (for k-nearest neighbors) in our approach. This systemic approach allows us to investigate both the existence and the hierarchical organization of conceptual clusters, directly addressing our research questions. Quantitative evaluation on external datasets (named entities 43, 12, numerical tokens, and social structures) demonstrates that token embeddings exhibit significant categorical community structure aligned with real-world concepts.
The structure of this paper is as follows: We first establish necessary background on embeddings, semantic representations, and evaluation strategies (Section 2), followed by a description of our methodological approach (Section 3). We then present our core findings, starting with LLM-human alignment (Section 4), with a focus on within-cluster properties and hierarchical structure. Then, we explore LLM-LLM alignment within their input embedding space (Section 5). Section 6 demonstrates the practical implications of our work through embedding engineering and bias mitigation. Finally, we conclude in Section 7.
2 Preliminaries
2.1 Static, Contextual and Base Embeddings
In this section, we clarify the distinctions between static, contextual, and base embeddings, which are crucial for understanding modern language models.
  • Static embeddings (e.g., Word2Vec33, GloVe 40) are context-independent vector representations of words, meaning each word has a fixed embedding regardless of its surrounding text. This limits their ability to handle polysemy (words with multiple meanings). These embeddings are typically pre-trained on large corpora and can be used in various downstream tasks. Critically, for the context of this work, static embeddings are product of legacy LM models and not used as the input representations of modern transformer-based LLMs. They are not inputs to transformer blocks and thus have limited significance when it comes to applications such as mitigation techniques (e.g. embedding engineering) in LLMs.
  • Contextual embeddings These are dynamic, context-dependent vectors. The embedding of a token is a function of its surrounding text, enabling the representation of nuanced meaning and resolving polysemy (e.g., Bert 11, Albert 23, GPT variants 42, 1). Different model layers capture different levels of contextual conditioning. However, this context-dependence limits direct interpretability and generalizability of individual token embeddings outside of specific contexts.
  • Base Embeddings: The process of generating contextual embeddings starts with base embeddings, which provide the initial vector representation for each input token. These differ from static embeddings such as GloVe and Word2Vec from the following perspectives. (1) Generation: Base embeddings are learned parameters within the LLM, as compared to separately trained static embeddings. (2) Usage: They are the direct input to the transformer blocks, forming the basis upon which contextualized representations are built through the model’s subsequent layers. These embeddings are the focus of our study.
2.2 Previous Works on Embedding Interpretability
Previous research on interpretability in LLMs has primarily focused on analyzing either contextual embeddings (for modern LLMs) or static embeddings (for legacy language models which are not directly applicable to LLMs due to the architectural difference).
In the realm of contextual representations, initial research focused on the learning dynamics of linguistic features within LLMs 50, 27, the scope has expanded to explore how these models acquire and represent knowledge about the world. Some studies 38, 19, 2 have investigated the ability of transformer-based models to learn representations of color, spatial, and temporal information. These studies often rely on analyzing contextual embeddings, which are intermediate outputs of LLMs. Moreover, contextual embeddings are inherently tied to the specific input context, making them highly variable and less generalizable.
The closest work to our knowledge that focuses on context agnostic embeddings within the modern LLMs is 7 where the authors propose a method to create context agnostic word embeddings from contextualized word representations using (sub)-word pooling as well as context combination techniques, and tested on semantic similarity datasets. Furthermore, 25 proposes a method for creating context-agnostic word representations by averaging the contextual embeddings derived from BERT, given a set of inferences on a masked target token within a corpus. In this methodology, the context is seen as a form of Gaussian noise that can be averaged out and hence produces a context-agnostic semantic representation. They observed such embedding represent richer semantic information than static word embedding counterparts (Word2Vec and GloVe) in intrinsic evaluation tasks. However, these works lack rigorous and extensive analysis of base embeddings of the LLMs to explores the intrinsic semantic organization within input embeddings of LLMs, possbile conceptual groupings and their alignments.
@@ 3 Concept Extraction
To investigate RQ1 and RQ2, we study the Human-LLM alignment of the input representations. To this end, we develop a method to first extract possible formed concepts within that space (refer to appendix A.1 for the discussion of conceptual groupings and semantic memory), and then to evaluate them against external datasets.
Our methodology consists of building the semantic graph, then using community detection to extract possible conceptual groupings. [^1] Note that we favor community detection over traditional clustering for identifying the conceptual clusters due to (1) its ability to align with the network-like structure of semantic representations, (2) its independence from the need for a predetermined number of clusters, and (3) its effectiveness in managing high-dimensional data by transforming it into a graph. Note that differentiable embedding functions guarantee a smooth, optimizable embedding space, but do not ensure a uniform distribution of concept instances. This unevenness, arising from factors like varying concept complexity and instance frequency, implies that while conceptual groupings may be locally uniform, the overall embedding space can be unevenly distributed. Therefore, effectively identifying these groupings requires mitigating this unevenness, which we address using a UMAP-based fuzzy graph construction.
Data: All tokens in the input embeddings.
Result: A set of hierarchical communities.
1 Create a community list. The initial community is the entire space;
2 for k=[different neighbor sizes] do
3       for all communities do
4             - Generate knn graph from the input embedding weights;
5             - Compute the edge weights of the graph using fuzzy simplex;
6             - Apply Louvain community detection;
7             - Add the identified communities to the list;
8
9
Algorithm 1 Concept Extraction
@@ 3.1 Graph Construction
The first step in the Uniform Manifold Approximation and Projection (UMAP) algorithm is to approximate the manifold by constructing a fuzzy topological representation of the embeddings using nearest neighbor descent 32. Inspired by that, we use the same nearest neighbor descent method to find the $K$ nearest neighbors for every token embedding in the embedding space and then use the same equations used in UMAP’s fuzzy graph construction to define the weight function of the edge between $x_{i}$ and $x_{j}$ nodes in the $K$-NN graph 32:
@@
Building upon our theoretical arguments, the conceptual/categorical representations (if they exist), should form fuzzy partitions that can be detected by the community detection algorithms. Note that since UMAP dimensionality reduction process can lead to information loss 15, 56, potentially obscuring important nuances in the representations, we perform community detection in the high-dimensional space.
@@ 3.2 Louvain Community Detection
The Louvain method is a widely used algorithm for community detection in large networks. It finds the communities by optimizing a metric called modularity. The modularity of a partition is a scalar value between -1 and 1 that measures how much more densely connected the nodes within a community are compared to how connected they would be in a random network. 6. For a weighted graph, modularity is defined as: